# The Griffiths phase and beyond: A large deviations study of the magnetic susceptibility of the two-dimensional bond-diluted Ising model

### v1

@@ IV Results for the disordered phase and the critical point
$\displaystyle x=\begin{cases}x_{-}=\dfrac{\chi-\overline{\chi}}{\overline{\chi}-1}&\text{if}~{}\chi<\overline{\chi},\\[8.53581pt] x_{+}=\dfrac{\chi-\overline{\chi}}{\chi_{\mathrm{f}}-\overline{\chi}}&\text{if}~{}\chi\geq\overline{\chi},\\ \end{cases}$ (17)
@@
Finally, Fig. 10 shows the distribution of $\chi$ at the critical temperature. The distribution has a concave shape, it is very flat and covers the whole range of the support from $\chi=1$ to $\chi_{f}\sim N$. The mean of the distribution diverges according to $\overline{\chi}\sim L^{\gamma_{f}/\nu_{f}}$, where $\gamma_{f}=7/4$ and $\nu_{f}=1$ are the critical exponents of the pure ferromagnet [10, 9]. To the left of the mean there is no convergence to a rate function. Instead a good data collapse is obtained by only rescaling the $x$-axis according to $(\chi-\overline{\chi})/\overline{\chi}$, without taking the log of the probabilities and rescaling. The origin of this behavior is the lack of self averaging at criticality in case of the diluted ferromagnet [50, 51]. To the right of the mean the large-deviation principle seems to be satisfied and one obtains a good collapse of the data by using a scaling rule according to Eq. (17).
@@ V Results for the ferromagnetic phase and at zero temperature
Figure 13 shows the distribution of the magnetic susceptibility at zero temperature for multiple system sizes. Again, it is visible that the tail of the distribution becomes wider on increasing the system size. Note that in this case the large-deviation approach even allows us to sample over the full support of the distribution. The support extends from the minimal possible value $\chi=0$, when there is only one cluster which contains all spins, to the maximum value $\chi=N/4$, when there are two clusters of equal size which together contain all spins. This is a direct consequence of Eq. (13) if one approaches the infinite cluster by considering the largest cluster of the finite-size systems under consideration. Figure 13 contains a heat map which illustrates the relation between the magnetic susceptibility and the local fraction of ferromagnetic bonds. It shows that the fraction of ferromagnetic bonds remains almost constant over a wide range of the magnetic susceptibility. Only for large values of $\chi$ there is a notable increase in the fraction. In comparison to Figure 12, which shows a similar heat map at $T=1.5$ inside the Griffiths phase, the interval in which $p_{\bm{J}}$ varies is much smaller. Instead the density of the second largest cluster $\rho_{2}$ is the driving force for the values of the magnetic susceptibility. The inset of Fig. 13 shows a heat map of the conditional probability $P(\rho_{2}^{2}|\chi)$. It demonstrates that there exists a linear relation between the square of the density of the second largest cluster and the magnetic susceptibility. This central importance of the second largest cluster can also be directly visualized by looking at bond samples with distinct different values of $\chi$ as they are shown in Fig. 14. The figure illustrates that for bond samples with a magnetic susceptibility which is close to the mean value the size of the second largest cluster is not important. For bond samples which originate from the plateau region of the distribution, the second largest cluster becomes significant, and for the extreme tail events there are only two clusters left with nearly identical size.
@@
Another interesting fact is that the mean of the zero-temperature distribution is non-zero, $\overline{\chi}=0.16016(5)$. This implies that a field-driven phase transition emerges for very low temperatures since $\partial\overline{m}/\partial h=\overline{\chi}/T$ such that $\partial\overline{m}/\partial h\sim T^{-1}$ for $T\to 0$. The external magnetic field destroys the ground state degeneracy because all spins will align in parallel to the magnetic field.
@@ VI Discussion
The distribution of the magnetic susceptibility is also studied directly at the critical temperature and inside the ferromagnetic phase. At the critical temperature a lack of self-averaging is observed to the left of the distribution mean, i.e., for small values of $\chi$ [50, 51]. Inside the ferromagnetic phase the distribution of the magnetic susceptibility exhibits an exponential tail similar to that of the Griffiths phase. The tail becomes wider with increasing system size and it is expected to extend to infinity in the thermodynamic limit. It is found that the driving force behind large values of the magnetic susceptibility in the ordered phase is a combination of the variance of the largest FKCK cluster and the size of the second largest cluster. At zero-temperature the second term, the size of the largest cluster of ferromagnetic bonds, is the only contribution to the large susceptibility observed.
@@
The Griffith phase is not particular to the diluted ferromagnet, but it can also be observed in other disordered systems, and we hope that the present work will motivate similar studies of related phenomena in related models. Of particular interest could be the case of the two-dimensional Ising spin glass [20]. This model exhibits frustrated interactions and does not have a ferromagnetic phase transition. Thus it would be interesting to see how the distribution of the magnetic susceptibility is impacted by these differences. Furthermore, we expect that similar effects of broad distributions will also be visible in observables other than the magnetic susceptibility. For the diluted ferromagnet studied here, for instance, the distribution of the specific heat would be of special significance since the specific heat describes the fluctuations of the internal energy that becomes singular at the ferromagnetic phase transition just as the magnetic fluctuations represented in the susceptibility. Finally, it would me most intriguing to apply the rare-event sampling techniques showcased here also for the case of the Griffiths singularities observed in quantum systems [21, 22, 23, 24, 25, 26, 27].
@@ Appendix B Exponential tail according to Bray
Below $T_{\mathrm{f}}$ and outside of the critical region, we have $\chi_{\mathrm{f}}\sim N$ such that one can define the intensive quantity $x=\chi/N$ which gives $P_{x}(x;N)=\exp(-AxN-2\ln(x)-\ln(N))$. The corresponding rate function then is given by

### v2

@@ IV Results for the disordered phase and the critical point
$\displaystyle x=\begin{cases}x_{-}=\dfrac{\chi-\overline{\chi}}{\overline{\chi}-1}&\text{if}~{}\chi<\overline{\chi}\\[8.53581pt] x_{+}=\dfrac{\chi-\overline{\chi}}{\chi_{\mathrm{f}}-\overline{\chi}}&\text{if}~{}\chi\geq\overline{\chi}\\ \end{cases}$ (17)
@@
Finally, Fig. 10 shows the distribution of $\chi$ at the critical temperature. The distribution has a concave shape, it is very flat and covers the whole range of the support from $\chi=1$ to $\chi_{f}\sim N$. The mean of the distribution diverges according to $\overline{\chi}\propto L^{\gamma_{f}/\nu_{f}}[1+\mathcal{O}(1/\ln L)]$, where $\gamma_{f}=7/4$ and $\nu_{f}=1$ are the critical exponents of the pure ferromagnet and $\mathcal{O}(1/\ln L)$ refers to the “big $\mathcal{O}$ notation” [16]. To the left of the mean there is no convergence to a rate function. Instead a good data collapse is obtained by only rescaling the $x$-axis according to $(\chi-\overline{\chi})/\overline{\chi}$. The origin of this behavior is the lack of self averaging at criticality in case of the diluted ferromagnet [60, 61, 62]. To the right of the mean the large-deviation principle seems to be satisfied and one obtains a good collapse of the data by using a scaling rule according to Eq. (17).
@@ V Results for the ferromagnetic phase and at zero temperature
Figure 13 shows the distribution of the magnetic susceptibility at zero temperature for multiple system sizes. Again, it is visible that the tail of the distribution becomes wider on increasing the system size. Note that in this case the large-deviation approach even allows us to sample over the full support of the distribution. The support extends from the minimal possible value $\chi=0$, when there is only one cluster which contains all spins, to the maximum value $\chi=N/4$, when there are two clusters of equal size which together contain all spins. This is a direct consequence of Eq. (13) if one approaches the infinite cluster by considering the largest cluster of the finite-size systems under consideration. Figure 13 contains a heat map which illustrates the relation between the magnetic susceptibility and the local fraction of ferromagnetic bonds. It shows that the fraction of ferromagnetic bonds remains almost constant over a wide range of the magnetic susceptibility. Only for large values of $\chi$ there is a notable increase in $p_{\bm{J}}$. In comparison to Figure 12, which shows a similar heat map at $T=1.5$ inside the Griffiths phase, the interval in which $p_{\bm{J}}$ varies is much smaller. Instead the density of the second largest cluster $\rho_{2}$ is the driving force for the values of the magnetic susceptibility. The inset of Fig. 13 shows a heat map of the conditional probability $P(\rho_{2}^{2}|\chi)$. It demonstrates that there exists a linear relation between the square of the density of the second largest cluster and the magnetic susceptibility. This central importance of the second largest cluster can also be directly visualized by looking at bond samples with distinct different values of $\chi$ as they are shown in Fig. 14. The figure illustrates that for bond samples with a magnetic susceptibility which is close to the mean value the size of the second largest cluster is not important. For bond samples which originate from the plateau region of the distribution, the second largest cluster becomes significant, and for the extreme tail events there are only two clusters left with nearly identical size.
@@
Another interesting fact is that the mean of the zero-temperature distribution is non-zero, $\overline{\chi}=0.16016(5)$. We interpret this as a field-driven continuous phase transition that emerges at very low temperatures because the finite-size clusters will align parallel to the magnetic field, i. e. $\partial\overline{m}/\partial h=\overline{\chi}/T$ such that $\partial\overline{m}/\partial h\sim T^{-1}$ for $T\to 0$. The external magnetic field destroys the ground state degeneracy.
@@ VI Discussion
The distribution of the magnetic susceptibility is also studied directly at the critical temperature and inside the ferromagnetic phase. At the critical temperature a lack of self-averaging is observed to the left of the distribution mean, i.e., for small values of $\chi$ [60, 61, 62]. Inside the ferromagnetic phase the distribution of the magnetic susceptibility exhibits an exponential tail similar to that of the Griffiths phase. The tail becomes wider with increasing system size and it is expected to extend to infinity in the thermodynamic limit. It is found that the driving force behind large values of the magnetic susceptibility in the ordered phase is a combination of the variance of the largest FKCK cluster and the second moment of the second largest cluster. At zero-temperature the size of the second largest cluster of ferromagnetic bonds is the only contribution to the large susceptibility observed.
@@
In the present work, we have focused on the distribution of the susceptibility for a fixed expectation value of the fraction of ferromagnetic bonds, namely $p=0.6$. This value of $p$ was chosen because it allows for a large Griffiths phase but is still far away from the percolation threshold $p_{\text{th}}=0.5$ such that there exist a low temperature ferromagnetic phase. In general we expect our results to be independent of the value of $p$ in such a sense that an exponential tail should exist for all values of $T$ and $p$ inside the Griffiths phase as well as in the ferromagnetic phase, but its shape may look different, see Appendix D for more details. Furthermore we note that for $T=0$ the magnetic susceptibility is equal to the average cluster size of the ferromagnetic bond clusters. The problem thereby simplifies to random-bond percolation, suggesting that analytical progress could potentially be made in order to derive the rate function.
The Griffiths phase is not particular to the diluted ferromagnet, but it can also be observed in other disordered systems, and we hope that the present work will motivate similar studies of related phenomena in other models. Of particular interest could be the case of the two-dimensional Ising spin glass [30]. This model exhibits frustrated interactions and does not have a ferromagnetic phase transition. Thus it would be interesting to see how the distribution of the magnetic susceptibility is impacted by these differences. Furthermore, we expect that similar effects of broad distributions will also be visible in observables other than the magnetic susceptibility. For the diluted ferromagnet studied here, for instance, the distribution of the specific heat would be of special significance since the specific heat describes the fluctuations of the internal energy that becomes singular at the ferromagnetic phase transition just as the magnetic fluctuations represented in the susceptibility [67]. Finally, it would me most intriguing to apply the rare-event sampling techniques showcased here also for the case of the Griffiths singularities observed in quantum systems [31, 32, 33, 34, 35, 36, 37].
@@ Appendix B Exponential tail according to Bray
Below $T_{\mathrm{f}}$ and outside of the critical region, we have $\chi_{\mathrm{f}}\sim N$ such that one can define the intensive quantity $x=\chi/N$ which gives $P_{x}(x;N)\propto\exp(-AxN-2\ln(x)-\ln(N))$. The corresponding rate function then is given by
@@ Appendix C Ferromagnetic bond clusters
Figure 18: Distribution of the magnetic susceptibility for $p=0.8$ and $T=2$ (inside the Griffiths phase) for different system sizes.
Figure 19: The main plot shows the rate function of the magnetic susceptibility for $p=0.8$ and $T=2$ right of the mean and the inset (a) illustrates the rate function left of the mean. The inset (b) depicts the rate function right of the mean for $p=0.6$ and $T=2$. The $x$-axis is rescaled according to Eq. (17).
Appendix D Influence of the occupation probability $p$
Here we discuss the potential effects of varying the bond occupation probability $p$. The arguments of Bray [24] which explain the existence of an exponential tail in the distribution of magnetic susceptibility inside the Griffiths phase do not depend on a specific value of $p$, but they are generally applicable for $10<p<$ 1. The shape of the distribution, however, can be different for different values of $p$. Figure 9 demonstrates the temperature dependence of the shape of the distribution. In a similar way we expect the distribution to look different when the value of $p$ is changed, but the presence of an exponential tail inside the Griffiths phase, in general, should not be affected by the specific value of $p$. As an example Fig. 18 illustrates the distribution for $p=0.8$ and $T=2$, inside the Griffiths phase. As expected, the mean value of $\chi$ is larger than that of $p=0.6$ at the identical temperature, cf. Fig. 6, but the exponential tail is again clearly visible. Furthermore, Fig. 19 depicts the rate function of the distribution and demonstrates that the range of the tail increases with system size. This rate function exhibits the same characteristic shape as the rate function for $p=0.6$ at $T=2$, as shown in the inset (b) for comparison, but it looks more similar to that at temperature $T=1.5$ for $p=0.6$, see Fig. 7. We have also performed test simulations for $p=0.3$ and again observed an exponential tail in agreement to the theoretical predictions [24]. Moreover we have performed a simulation for $T=0$ when $p=0.8$ and found a similar behavior as described in Sec. V. We hence find broad numerical confirmation of the expectation that the existence and general featuers of theb Griffiths phase are independent of the bond occupation probability $p$.